# Solar flares in the Solar Orbiter Era: Short exposure EUI/FSI observations of STIX flares

### v1

@@ 3.2 STX2022-11-13T06:18: A “standard” flare
STX2022-11-13T06:18 was a C1.4 GOES-class flare observed by both Solar Orbiter and Earth based observatories. At this time Solar Orbiter was at a distance of 0.64 AU from the Sun and an angle of to the Sun-Earth line. Two EUI/FSI short exposure frames were obtained during the flare, one during the first non-thermal HXR burst and the second towards the end of the flare (see Fig. 3 a). Figures 3 b) and c) are composite maps of the normal and short exposure EUI/FSI frames from the times denoted by the vertical dashed lines in panel a). In such composite maps, all non-zero short exposure pixel values have replaced their corresponding normal exposure pixel counts. The counts were normalised by exposure time and the flux correction factor was applied as for STX2023-07-16T04:32. The corresponding AIA 171 Å observations are shown in Fig. 6.
@@
This flare is another example of a flare whose signatures are consistent with the standard flare ‘CSHKP’ model (Carmichael, 1964; Sturrock, 1966; Hirayama, 1974; Kopp & Pneuman, 1976). During the first HXR burst we see non-thermal emission at 15-28 keV from two footpoints and thermal emission from the base of the flare loop emitted at 6-10 keV (Fig. 3 b). The time interval used to reconstruct the STIX images in Fig. 3 b) was 06:15:30-06:16:15 UT. Interestingly two distinct brightenings are seen in EUI/FSI 174 Å which align well to the HXR footpoints. Later in the flare, the previously heated thermal loop connecting the two footpoints is observable in the 174 Å EUI/FSI channel after cooling to temperatures the channel is the most sensitive to (the temperature response peaks at 1 MK). The EUI short exposure frame shows a low-lying or tilted loop formed in the early phase of the flare and extending between the two footpoints observed in the previous frame. In contrast, the 6-10 keV emission (for the interval 06:23:05-06:23:35 UT) shows the hotter, higher and more arch-shaped loops formed during the later phase of the flare.
@@
This example demonstrates the need for high cadence short exposure observations. The EUI/FSI channels are sensitive to a range of temperatures and therefore see both the impulsive impact of non-thermal particles as well as more gradual heating in flares. The short exposure EUI/FSI frame enables us to identify the brightest of the saturated pixels and hence enables us to confidently align the STIX non-thermal sources. Fortunately in this case, there was a short exposure frame which captured the heating of flare ribbons caused by the impact of non-thermal particles as well as post flare heating/cooling, However, since the response of the 174 Å channel is temperature sensitive and thus changes significantly in time, in order to investigate the exact correspondence between EUV and HXR emission in flares, high time cadence observations are required.
@@
In order to verify our understanding of how the emission observed in the 174 Å passband short exposure flare observations relates to the injected electron beam, we simulate the atmospheric response to an injected electron beam, the parameters of which are inferred from HXR measurement. Furthermore, as this flare is a classic example of the standard flare model at work, it lends itself well to modelling efforts. We do this using the 1D radiative hydrodynamic RADYN code (Carlsson & Stein, 1992, 1995, 1997, 2002; Abbett & Hawley, 1999; Allred et al., 2005, 2015). The input electron beam parameters were derived by fitting the HXR spectrum during the first non-thermal burst (06:15:29-06:16:16 UT) with a thermal model (f_vth.pro), a thick target bremsstrahlung component (f_thick2.pro) and an albedo component using the OSPEX spectral fitting software. The data between 4-36 keV was included in the fitting procedure and the resultant fit obtained a reduced chi-square, , assuming 5% systematic errors. The fitted models indicated a non-thermal electron beam with electron rate, electrons/s, a spectral index, and low energy cutoff, keV, from which we obtain a total non-thermal power of erg/s. The power obtained is an estimate of the total spatially integrated power in the electron beam. In order to estimate the area over which this electron beam is distributed we calculate the footpoint area. In this case the advantage of having unsaturated footpoint observations was weakened by the fact that the spatial resolution of EUI/FSI at this radial distance is lower than that of AIA. Therefore estimating the footpoint area using the EUI/FSI short exposure frame gives a similar result to AIA. We estimated the footpoint area by calculating the area of the pixels within the 30% AIA contour level during the first non-thermal burst. The choice of the contour level is somewhat arbitrary, however in this case 30% was chosen as a middle ground between flux over- and underestimation. From this we obtain a total footpoint area estimate of . This results in a total electron beam energy flux of erg/s/ which is used as input into the RADYN model. Furthermore, the loop length was estimated by simply using half the distance between footpoint brightenings in AIA 171 Å as an estimate of the radius of a semi-circular loop, from which a half-loop length (from loop apex to footpoint) of approximately 17 Mm was obtained.
@@
An important consideration is the impact of the choice of starting atmosphere on the model results. In the RADYN run presented here, a starting apex temperature of 3 MK and a simple starting atmosphere (VAL3C), similar to what is used in the FCHROMA grid database (Carlsson et al., 2023) was used. Polito et al. (2018) demonstrated the impact of initial loop-top temperatures for nanoflare emission. In particular, Polito et al. (2018) showed that for higher starting temperatures, electrons are typically stopped at high altitudes in the corona and therefore only a small amount of energy is dissipated in the transition region and chromosphere. This has a direct impact of the intensity of EUV emission.
@@
A triangular beam of duration 45 s with the aforementioned energy flux, low energy cutoff of 13 keV and spectral index of 5 was input in the model and the 1D atmospheric response simulated. The model outputs include a grid of temperature and ambient ionised hydrogen density as a function of height along the loop as a function of time. The emission measure and temperature in each grid cell was then folded through the latest EUI/FSI 174 Å response (2024, priv. comm. with Frédéric Auchere) in order to determine the predicted response. Since this wavelength range is optically thin, the total response can be obtained by integrating along the entire loop length.
@@
It is important to note that the simulated grid is adaptive, which means that the cell boundaries vary in time in order to enable accurate simulation of shocks and sharp discontinuities. For representation purposes, Fig. 4 b) is obtained by interpolating the temperature and density grids to a fine (0.0003 Mm resolution), evenly sampled grid and folding the resampled grid through the EUI/FSI 174 Å response function. This introduces minor errors due to interpolation, however, for the purpose of demonstrating the height of the observed emission, the interpolation accuracy is sufficient.
@@
Fig. 4 a), shows the predicted EUI/174 Å flux given the simulated atmosphere, integrated in height. The EUI/FSI short exposure frame was taken approximately 16.5 s into the beam injection, at this time the total predicted EUI/FSI 174 Å flux is DN/s/pix which agrees remarkably well with the observed peak flux in the short exposure frame whose maximum is DN/s/pix. Fig. 4 b), shows the predicted 174 Å flux as a function of height and time. As expected, the majority of the flux detected originates from a very thin layer in the chromosphere (0.9-1.3 Mm above the photosphere). This indicates that the majority of the 174 Å EUV flux during the impulsive phase of the flare originated from the chromosphere. In comparison, coronal emission is very faint due to both the significantly higher temperatures and lower emission measure at this time.
@@ 3.3 STX2023-04-22T22:21: Potential anchor points of an erupting filament
STX2023-04-22T22:21 was an estimated M1 GOES-class flare (Xiao et al., 2023). It occurred while Solar Orbiter was at a distance of 0.4 AU from the Sun and an angle of from the Sun-Earth line. The flare was not observed by Earth based instruments. Fig. 5 a) shows the STIX lightcurves with the time of the short exposure observations denoted by the vertical dashed lines. Panels b) and c) show the EUI/FSI short exposure observations at these times with STIX contours overlaid. The blue contours in b) are the 5-85% contours levels of the 18-50 keV emission during this time. The map has been created by forward-fitting four circular Gaussian sources with all other parameters (location, source size and flux) free for the time range 22:20:30-22:21:10 UT, using the method detailed by Volpara et al. (2022). The position and error on the coordinates are shown. The red contours in panels b) and c) denote the 20-80% contours levels of the 6-10 keV emission of both 40 s intervals centred around the EUI/FSI short exposure time. In panel c), there is emission from previously heated loops which have presumably cooled to temperatures the EUV passband is highly sensitive to. The thermal HXR emission at this time originates from an elongated source which spans the extent of the original non-thermal structure shown in panel b).
The four non-thermal sources of HXR emission are of interest here as they are thought to correspond to the anchor points of an erupting flux rope. Stiefel et al. (2023) recently observed non-thermal HXR emission from the footpoints of an erupting filament alongside the standard flare footpoint emission. Similar observations have been reported by Chen et al. (2020) in microwave wavelengths, who observed microwave emission coming from the filament legs during the impulsive phase of the flare. In both cases, the emission was interpreted as a signature of non-thermal electrons which propagate along the flux rope post reconnection and finally reach chromospheric altitudes resulting in the observed HXR and microwave emission respectively. Such observations can be used to study transport effects on particles in an erupting flux rope. Furthermore, such observations can shed light on the acceleration process of electrons in flares.
As Earth observations of this flare are not available, EUI/FSI observations are essential for providing context to STIX observations. This is particularly important here as four sources are resolved with STIX, the faintest of which is only at the 5% contour level. This observation is therefore pushing the boundary of what STIX is capable of resolving given its limited dynamic range. However, given the strong correspondence with the brightest EUV sources seen in the short exposure frame, we are confident that STIX is reliably resolving four real non-thermal sources.
@@ 4 Discussion and summary
4.1 The importance of the EUI/FSI short exposure observing mode
@@
Short exposure EUV/UV observations are a key observational asset for flare studies and are particularly important for missions like Solar Orbiter, that frequently observe the Sun from the “back-side” and therefore do not have context observations from Earth based assets. Short exposure observations of flares are important for precisely identifying flare locations. Specifically, EUI/FSI short exposure observations can be utilised to correct the absolute pointing of STIX which is a significant advantage of having the EUI and STIX instruments on the same spacecraft. These observations are also useful for identifying fine structure within flaring ribbons, active regions and other transients such as coronal mass ejections. EUI/FSI onboard Solar Orbiter has been routinely collecting such data since November 2022, with more than 9000 flares co-observed by STIX and EUI/FSI from then until December 2023.
4.2 Extending the short exposure mode to EUI/HRIEUV
@@
While EUI/FSI observations are critical, what is really needed in order to advance our understanding of particle acceleration, energy release and transport in flares is high cadence short exposure EUV observations, with high spatial resolution. Very high-cadence (0.3-1.7s) images and spectral observations from the Interface Region Imaging Spectrograph (IRIS; De Pontieu et al., 2014, 2021) have provided significant new insights on plasma dynamics occurring in flares, that cannot be observed at lower cadence (Jeffrey et al., 2018; French et al., 2021; Lörinčík et al., 2022). Furthermore, with high cadence, short exposure observations the spatial origin of rapid variations in the observed emission could be determined, a pursuit which is particularly important for the study of quasi-periodic pulsations in flares (e.g. Zimovets et al., 2021). Resolving the spatial origin of such variation is a key component in understanding particle acceleration in flares (e.g. Collier et al., 2024), see also discussion in Inglis et al. (2023). These recent studies have motivated the need for regular, unsaturated and high-cadence flare observations.
@@
In order to do this, coordinated efforts are under way in which Solar Orbiter instruments are operating with flare optimised settings during remote sensing windows, with support from Earth based assets. In such modes, the suite of relevant remote sensing instruments onboard Solar Orbiter are pointed at the active region predicted to be the most likely to flare. During the recent perihelion of March & April 2024 many C-and several M- GOES class flares were observed during this flare campaign with completely unprecedented resolution. For example, the M-class flare of March 19th 2024 was observed while EUI/HRIEUV was operating at high cadence (2 s), taking both normal (2 s) and short (0.04 s) exposure observations with a 2-pixel resolution of 310 km. Such observations promise enlightening insights into particle acceleration, heating and energy partition in solar flares. A follow-on paper will discuss the high-cadence, short exposure EUI/HRIEUV data obtained during the March/April 2024 flare campaign in more detail. Another major flare campaign is planned for the perihelion of Solar Orbiter in March/April 2025.
@@ 4.3 Potential new science
Short exposure EUV/UV observations provide complementary diagnostics of the flaring region. “Far-side” context observations are becoming increasingly pertinent in the era of multi-messenger solar physics. In this work, we have highlighted the invaluable data that the EUI/FSI instrument onboard Solar Orbiter is now routinely obtaining. These new datasets reveal fine structure in flaring ribbons which is typically lost due to instrument saturation.
Importantly, from these observations, better constraints on the energy flux of non-thermal electron beams, required as input for flare modelling can be derived, particularly when Solar Orbiter is at a close radial distance to the Sun. This will be especially revolutionary given the unprecedented spatial resolution of EUI/HRIEUV (200 km 2-pixel resolution on the surface of the Sun) at perihelion (Berghmans et al., 2023). By combining unsaturated high cadence observations of flaring ribbons from HRIEUV and STIX onboard Solar Orbiter, the non-thermal electron energy flux in flares can be determined much more precisely than previously possible. Finally, high cadence unsaturated observations of flaring loops will enable us to study flare heating and cooling processes in more detail.
In summary, the need for unsaturated EUV/UV observations of solar flares has been demonstrated in this paper. Currently, EUI onboard Solar Orbiter is providing excellent observations of this kind. Looking to the future, we note that there is a strong need for a high-cadence flare focused EUV/UV imager for the next solar cycle (Inglis et al., 2023).

### v2

@@ 3.2 STX2022-11-13T06:18: A ‘standard’ flare
STX2022-11-13T06:18 was a C1.4 GOES-class flare observed by both Solar Orbiter and Earth-based observatories. At this time, Solar Orbiter was at a distance of 0.64 AU from the Sun and an angle of to the Sun-Earth line. Two EUI/FSI short-exposure frames were obtained during the flare, one during the first non-thermal HXR burst and the second towards the end of the flare (see Fig. 3 a). Figures 3 b) and c) are composite maps of the normal- and short-exposure EUI/FSI frames from the times denoted by the vertical dashed lines in panel a). In such composite maps, all non-zero short-exposure pixel values have replaced their corresponding normal exposure pixel counts. The counts were normalised by exposure time and the flux correction factor was applied as for STX2023-07-16T04:32. The corresponding AIA 171 Å observations are shown in Fig. 6.
@@
This flare is another example of a flare whose signatures are consistent with the standard flare ‘CSHKP’ model (Carmichael, 1964; Sturrock, 1966; Hirayama, 1974; Kopp & Pneuman, 1976). During the first HXR burst, we see non-thermal emission at 15-28 keV from two footpoints and thermal emission from the base of the flare loop emitted at 6-10 keV (Fig. 3 b). The time interval used to reconstruct the STIX images in Fig. 3 b) was 06:15:30-06:16:15 UT. Interestingly, two distinct brightenings are seen in EUI/FSI 174 Å , which align well with the HXR footpoints. Later in the flare, the previously heated thermal loop connecting the two footpoints is observable in the 174 Å EUI/FSI channel after cooling to temperatures the channel is most sensitive to (the temperature response peaks at 1 MK). The EUI short-exposure frame shows a low-lying or tilted loop formed in the early phase of the flare and extending between the two footpoints observed in the previous frame. In contrast, the 6-10 keV emission (for the interval 06:23:05-06:23:35 UT) shows the hotter, higher, and more arch-shaped loops formed during the later phase of the flare.
@@
This example demonstrates the need for high-cadence short-exposure observations. The EUI/FSI channels are sensitive to a range of temperatures, and therefore see both the impulsive impact of non-thermal particles as well as more gradual heating in flares. The short-exposure EUI/FSI frame enables us to identify the brightest of the saturated pixels, and hence enables us to confidently align the STIX non-thermal sources. Fortunately, in this case there was a short-exposure frame that captured the heating of flare ribbons caused by the impact of non-thermal particles as well as post-flare heating or cooling. However, since the response of the 174 Å channel is temperature-sensitive, and thus changes significantly in time, in order to investigate the exact correspondence between EUV and HXR emission in flares, high-time-cadence observations are required.
@@
In order to verify our understanding of how the emission observed in the 174 Å passband short-exposure flare observations relates to the injected electron beam, we simulated the atmospheric response to an injected electron beam, the parameters of which are inferred from HXR measurement. Furthermore, as this flare is a classic example of the standard flare model at work, it lends itself well to modelling efforts. We did this using the 1D radiative hydrodynamic RADYN code (Carlsson & Stein, 1992, 1995, 1997, 2002; Abbett & Hawley, 1999; Allred et al., 2005, 2015). The input electron beam parameters were derived by fitting the HXR spectrum during the first non-thermal burst (06:15:29-06:16:16 UT) with a thermal model (f_vth.pro), a thick target bremsstrahlung component (f_thick2.pro), and an albedo component using the OSPEX spectral fitting software. The data between 4-36 keV was included in the fitting procedure and the resultant fit obtained a reduced chi-square, , assuming 5% systematic errors. The fitted models indicated a non-thermal electron beam with an electron rate of electrons/s, a spectral index of , and a low energy cut-off of keV, from which we obtained a total non-thermal power of erg/s. The power obtained is an estimate of the total spatially integrated power in the electron beam. In order to estimate the area over which this electron beam is distributed, we calculated the footpoint area. In this case, the advantage of having unsaturated footpoint observations was weakened by the fact that the spatial resolution of EUI/FSI at this radial distance is lower than that of AIA. Therefore, estimating the footpoint area using the EUI/FSI short-exposure frame gives a similar result to AIA. We estimated the footpoint area by calculating the area of the pixels within the 30% AIA contour level during the first non-thermal burst. The choice of the contour level is somewhat arbitrary; however, in this case 30% was chosen as a middle ground between flux over- and underestimation. From this, we obtained a total footpoint area estimate of . This results in a total electron beam energy flux of erg/s/, which is used as an input in the RADYN model. Furthermore, the loop length was estimated by simply using half the distance between footpoint brightenings in AIA 171 Å as an estimate of the radius of a semi-circular loop, from which a half-loop length (from loop apex to footpoint) of approximately 17 Mm was obtained.
@@
An important consideration is the impact of the choice of starting atmosphere on the model results. In the RADYN run presented here, a starting apex temperature of 3 MK and a simple starting atmosphere (VAL3C), similar to what is used in the FCHROMA grid database (Carlsson et al., 2023), was used. Polito et al. (2018) have demonstrated the impact of initial loop-top temperatures for nanoflare emission. In particular, Polito et al. (2018) show that for higher starting temperatures, electrons are typically stopped at high altitudes in the corona, and therefore only a small amount of energy is dissipated in the transition region and chromosphere. This has a direct impact on the intensity of EUV emission.
@@
A triangular beam of duration 45 s with the aforementioned energy flux, a low energy cut-off of 13 keV, and spectral index of 5 were input in the model and the 1D atmospheric response was simulated. The model outputs include a grid of temperature and ambient ionised hydrogen density as a function of height along the loop as a function of time. The emission measure and temperature in each grid cell was then folded through the latest EUI/FSI 174 Å response (2024, priv. comm. with Frédéric Auchere) in order to determine the predicted response. Since this wavelength range is optically thin, the total response can be obtained by integrating along the entire loop length.
@@
It is important to note that the simulated grid is adaptive, which means that the cell boundaries vary in time in order to enable accurate simulations of shocks and sharp discontinuities. For representation purposes, Fig. 4 b) was obtained by interpolating the temperature and density grids to a fine (0.0003 Mm resolution), evenly sampled grid and folding the resampled grid through the EUI/FSI 174 Å response function. This introduces minor errors due to interpolation; however, for the purpose of demonstrating the height of the observed emission, the interpolation accuracy is sufficient.
@@
Figure 4 a) shows the predicted EUI/174 Å flux given the simulated atmosphere, integrated in height. The EUI/FSI short-exposure frame was taken approximately 16.5 s into the beam injection. At this time, the total predicted EUI/FSI 174 Å flux is DN/s/pix, which agrees remarkably well with the observed peak flux in the short-exposure frame whose maximum is DN/s/pix. Figure 4 b) shows the predicted 174 Å flux as a function of height and time. As was expected, the majority of the flux detected originates from a very thin layer in the chromosphere (0.9-1.3 Mm above the photosphere). This indicates that the majority of the 174 Å EUV flux during the impulsive phase of the flare originated from the chromosphere. In comparison, coronal emission is very faint due to both the significantly higher temperatures and lower emission measure at this time.
@@ 3.3 STX2023-04-22T22:21: Potential anchor points of an erupting filament
STX2023-04-22T22:21 was an estimated M1 GOES-class flare (Xiao et al., 2023). It occurred while Solar Orbiter was at a distance of 0.4 AU from the Sun and an angle of from the Sun-Earth line. The flare was not observed by Earth-based instruments. Figure 5 a) shows the STIX light curves with the time of the short-exposure observations denoted by the vertical dashed lines. Panels b) and c) show the EUI/FSI short-exposure observations at these times with STIX contours overlaid. The blue contours in b) are the 5-85% contours levels of the 18-50 keV emission during this time. The map has been created by forward-fitting four circular Gaussian sources with all other parameters (location, source size, and flux) free for the time range 22:20:30-22:21:10 UT, using the method detailed by Volpara et al. (2022). The position and error on the coordinates are shown. The red contours in panels b) and c) denote the 20-80% contours levels of the 6-10 keV emission of both 40 s intervals centred around the EUI/FSI short-exposure time. In panel c), there is emission from previously heated loops that have presumably cooled to temperatures that the EUV passband is highly sensitive to. The thermal HXR emission at this time originates from an elongated source that spans the extent of the original non-thermal structure shown in panel b).
The four non-thermal sources of HXR emission are of interest here as they are thought to correspond to the anchor points of an erupting flux rope. Stiefel et al. (2023) recently observed non-thermal HXR emission from the footpoints of an erupting filament alongside the standard flare footpoint emission. Similar observations have been reported at microwave wavelengths by Chen et al. (2020), who observed microwave emission coming from the filament legs during the impulsive phase of the flare. In both cases, the emission was interpreted as a signature of non-thermal electrons that propagate along the flux rope post-reconnection and finally reach chromospheric altitudes, resulting in the observed HXR and microwave emission, respectively. Such observations can be used to study transport effects on particles in an erupting flux rope. Furthermore, such observations can shed light on the acceleration process of electrons in flares.
As Earth observations of this flare are not available, EUI/FSI observations are essential for providing context to STIX observations. This is particularly important here as four sources are resolved with STIX, the faintest of which is only at the 5% contour level. This observation is therefore pushing the boundary of what STIX is capable of resolving given its limited dynamic range. However, given the strong correspondence with the brightest EUV sources seen in the short-exposure frame, we are confident that STIX is reliably resolving four real non-thermal sources.
@@ 4 Discussion and summary
4.1 The importance of the EUI/FSI short-exposure observing mode
@@
Short-exposure EUV/UV observations are a key observational asset for flare studies and are particularly important for missions like Solar Orbiter that frequently observe the Sun from the ‘back-side’ and that therefore do not have context observations from Earth-based assets. Short-exposure observations of flares are important for precisely identifying flare locations. Specifically, EUI/FSI short-exposure observations can be utilised to correct the absolute pointing of STIX, which is a significant advantage of having the EUI and STIX instruments on the same spacecraft. These observations are also useful for identifying fine structure within flaring ribbons, active regions, and other transients such as coronal mass ejections. EUI/FSI on board Solar Orbiter has been routinely collecting such data since November 2022, with more than 9000 flares co-observed by STIX and EUI/FSI from then until December 2023.
4.2 Extending the short-exposure mode to EUI/HRIEUV
@@
While EUI/FSI observations are critical, what is really needed in order to advance our understanding of particle acceleration, energy release, and transport in flares is high-cadence short-exposure EUV observations with high spatial resolution. Very high-cadence (0.3-1.7s) images and spectral observations from the Interface Region Imaging Spectrograph (IRIS; De Pontieu et al., 2014, 2021) have provided significant new insights into plasma dynamics occurring in flares, which cannot be observed at lower cadence (Jeffrey et al., 2018; French et al., 2021; Lörinčík et al., 2022). Furthermore, with high-cadence, short-exposure observations the spatial origin of rapid variations in the observed emission could be determined, a pursuit that is particularly important for the study of quasi-periodic pulsations in flares (e.g. Zimovets et al., 2021). Resolving the spatial origin of such variation is a key component in understanding particle acceleration in flares ((e.g. Collier et al., 2024); see also discussion in Inglis et al. (2023)). These recent studies have motivated the need for regular, unsaturated, and high-cadence flare observations.
@@
In order to do this, co-ordinated efforts are under way in which Solar Orbiter instruments are operating with flare-optimised settings during remote sensing windows, with support from Earth-based assets. In such modes, the suite of relevant remote sensing instruments on board Solar Orbiter are pointed at the active region predicted to be the most likely to flare. During the recent perihelion of March and April 2024, many C-and several M- GOES class flares were observed during this flare campaign with completely unprecedented resolution. For example, the M-class flare of March 19 2024 was observed while EUI/HRIEUV was operating at high cadence (2 s), taking both normal- (2 s) and short- (0.04 s) exposure observations with a 2-pixel resolution of 310 km. Such observations promise enlightening insights into particle acceleration, heating, and energy partition in solar flares. A follow-on paper will discuss the high-cadence, short-exposure EUI/HRIEUV data obtained during the March-April 2024 flare campaign in more detail. Another major flare campaign is planned for the perihelion of Solar Orbiter in March-April 2025.
@@ 4.3 Potential new science
Short-exposure EUV/UV observations provide complementary diagnostics of the flaring region. ‘Far-side’ context observations are becoming increasingly pertinent in the era of multi-messenger solar physics. In this work, we have highlighted the invaluable data that the EUI/FSI instrument on board Solar Orbiter is now routinely obtaining. These new datasets reveal fine structure in flaring ribbons that is typically lost due to instrument saturation.
Importantly, from these observations, better constraints on the energy flux of non-thermal electron beams, required as input for flare modelling, can be derived, particularly when Solar Orbiter is at a close radial distance to the Sun. This will be especially revolutionary given the unprecedented spatial resolution of EUI/HRIEUV (200 km 2-pixel resolution on the surface of the Sun) at perihelion (Berghmans et al., 2023). By combining unsaturated high-cadence observations of flaring ribbons from HRIEUV and STIX on board Solar Orbiter, the non-thermal electron energy flux in flares can be determined much more precisely than was previously possible. Finally, high-cadence unsaturated observations of flaring loops will enable us to study flare heating and cooling processes in more detail.
In summary, the need for unsaturated EUV/UV observations of solar flares has been demonstrated in this paper. Currently, EUI on board Solar Orbiter is providing excellent observations of this kind. Looking to the future, we note that there is a strong need for a high-cadence flare-focused EUV/UV imager for the next solar cycle (Inglis et al., 2023).